OpenStudy (anonymous):

Find all values of m so that the function y = emx is a solution of the given differential equation. (Enter your answers as a comma-separated list.) y' + 3y = 0

OpenStudy (kainui):

Where are you having trouble?

OpenStudy (anonymous):

Idk how to start. Do I integrate both sides first?

OpenStudy (anonymous):

or do I integrate \[y=e ^{mx}\] and then plug it in in the equation \[y \prime + 3y = 0\] ?

OpenStudy (bradely):

y'+3y=0 (D+3)y=0 D+3=0 D=-3 y=e^-3x Source: http://www.mathskey.com/question2answer/

OpenStudy (bradely):

or dy/dx+3y=0 dy/dx=-3y dy/y = -3dx Integrate both sides lny =-3x y =e^-3x+c Source: http://www.mathskey.com/question2answer/

OpenStudy (unklerhaukus):

\[y = e^{mx}\]What is the derivative of this? \[y'=. . .%me^{mx} \]

OpenStudy (unklerhaukus):

then plug \(y\) and \(y'\) into \[y' + 3y = 0\] and solve for m

OpenStudy (unklerhaukus):

Hint: (factor, and assume \(e^{mx}\neq0\))